how many grams of ammonia is produced from 12 g of hydrogen
Answer #1


Molar mass of H2 = 2.016 g/mol

mass of H2 = 12 g
mol of H2 = (mass)/(molar mass)
= 12/2.016
= 5.952 mol


Balanced chemical equation is:
N2 + 3H2 ---> 2NH3


According to balanced equation
mol of NH3 formed = (2/3)* moles of H2
= (2/3)*5.952
= 3.968 mol


Molar mass of NH3,
MM = 1*MM(N) + 3*MM(H)
= 1*14.01 + 3*1.008
= 17.034 g/mol


mass of NH3 = number of mol * molar mass
= 3.968*17.03
= 67.6 g
Answer: 68 g